# Theory of Valley Splitting in Si/SiGe Spin-Qubits: Interplay of Strain, Resonances and Random Alloy Disorder

### v1

@@ V Conclusions and Outlook
We have developed a comprehensive theoretical model that describes a broad range of physical mechanisms determining the valley splitting in Si/SiGe heterostructures, namely alloy-disorder, strain and resonances between valley states across neighboring Brillouin zones. The numerical results are consistent with previously known results but also offer new insights extending the state of the art. With regard to applications, the key accomplishment of the model presented here is a comprehensive characterization of the long-period wiggle-well. This includes complex phenomena such as the shear strain-induced symmetry breaking, which unlocks the corresponding resonance mechanism, as well as a quantification of the statistical fluctuations of the valley splitting caused by alloy disorder. Our model predicts a significant deterministic enhancement of the valley splitting for the long-period wiggle-well. Because of the critical role of shear strain, we conclude that strain engineering [42, 56, 32] must be further advanced to enhance the design of future Si/SiGe spin-qubit devices.
In contrast to detailed atomistic models, the present model offers a comprehensive description of multiple physical phenomena at very low computational cost. Therefore, it is well suited to rapidly characterize the valley splittings of a vast number of different epitaxial designs. In a subsequent work, the present model will be employed for free-shape optimization of the epitaxial profile to further enhance the valley splitting beyond the limitations of the simple parametrization considered here.
@@ Acknowledgements.
This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany´s Excellence Strategy – The Berlin Mathematics Research Center MATH+ (EXC-2046/1, project ID: 390685689, project AA2–17). M. K. is grateful for valuable discussions with Mark Friesen, Merritt P. Losert and Lars R. Schreiber.
@@ Appendix A Derivation of the Multi-Valley Coupled Envelope Equation Model
and the confinement potential $U\left(\mathbf{r}\right)$ described in Eq. (4) of the main text. Here, $m_{0}$ is the vacuum electron mass and $V_{0}\left(\mathbf{r}\right)$ is the lattice-periodic potential of the crystal. The crystal potential is invariant under translations with lattice vectors $V_{0}\left(\mathbf{r}\right)\equiv V_{0}\left(\mathbf{r}+\mathbf{R}_{j}\right)$. We assume that $V_{0}$ describes solely the perfectly periodic Si crystal, whereas the effects of Ge in both the QW and the barrier are described by the heterostructure potential that is part of $U\left(\mathbf{r}\right)$, see Eq. (4). In this way, double-counting of Ge atoms is avoided and alloy disorder effects are separated from the idealized band structure computation.
@@ A.1 Lattice-Periodic Part
The lattice-periodic part of the coupled envelope equations is obtained from considering
@@ A.2 Heterostructure Part
In the following, the corresponding computation is carried out for the heterostructure part, which finally induces a coupling between the different valley-specific envelope wave functions. The ansatz (48) with (50) yields
@@
This expression is separated into the intervalley ($\nu\neq\nu^{\prime}$) and the intravalley ($\nu=\nu^{\prime}$) contribution. Since the Bloch factors vary on an atomistic scale, i.e., much faster than the confinement potential and the envelope wave function, the intravalley part can be separated as [66]
@@ A.3 Coupled Envelopes
In the case of a single band (conduction band only, band indices $n$, $n^{\prime}$ are suppressed in the following), two valleys at $\mathbf{k}_{\nu}=\pm\mathbf{k}_{0}$, symmetry factors $\alpha_{\nu}=1/\sqrt{2}$ and identical effective masses $m_{n,\mathbf{k}_{\nu}}=m$ and band edge energies $E_{n,\mathbf{k}_{\nu}}=E_{c}$, we obtain
@@ Appendix B Evaluation of Valley Splitting Components and Band Structure Coefficients
In the small strain limit, the reciprocal lattice vectors are written as $\mathbf{G}_{n}=\sum_{j}n_{j}\left(I-\varepsilon\right)\mathbf{b}_{j}$, where the $\mathbf{b}_{j}$, $j\in\left\{1,2,3\right\}$ are the primitive reciprocal lattice vectors of the unstrained fcc lattice [54]
@@
These selection rules greatly simplify the evaluation of the summation in $\Delta_{\mathrm{det}}^{l}$, which (after renaming of the remaining free summation index to $n$) yields the compact result
@@ C.1 Intervalley Coupling Parameter
where we have introduced the abbreviation
@@
$=\mathrm{e}^{-\frac{1}{4}\left|s\right|^{2}\Gamma-\frac{1}{4}\mathrm{Re}\left(s^{*2}C\right)},$
which is the characteristic function of a scalar complex-valued normal distribution with covariance $\Gamma$ and relation parameter (or pseudo-covariance) $C$. This is essentially the statement of the central limit theorem for complex-valued random variables. Hence

### v2

@@ V Conclusions and Outlook
We have developed a comprehensive theoretical model that describes a broad range of physical mechanisms determining the valley splitting in Si/SiGe heterostructures, namely alloy-disorder, strain and resonances between valley states across neighboring Brillouin zones. The numerical results are consistent with previously known results but also offer new insights extending the state of the art. With regard to applications, the key accomplishment of the model presented here is a comprehensive characterization of the long-period wiggle-well. This includes complex phenomena such as the shear strain-induced symmetry breaking, which unlocks the corresponding resonance mechanism, as well as a quantification of the statistical fluctuations of the valley splitting caused by alloy disorder. Our model predicts a significant deterministic enhancement of the valley splitting for the long-period wiggle-well. Due to the critical role of shear strain, we conclude that strain engineering [43, 58, 33] must be further advanced to enhance the design of Si/SiGe spin-qubits.
In contrast to detailed atomistic models, the present model offers a comprehensive description of multiple physical phenomena at very low computational cost. Therefore, it is well suited to rapidly characterize the valley splittings of a vast number of different epitaxial designs. In a subsequent work, the present model will be employed for free-shape optimization of the epitaxial profile to further enhance the valley splitting beyond the simple heuristics considered here.
@@ Acknowledgements.
This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – The Berlin Mathematics Research Center MATH+ (EXC-2046/1, project ID: 390685689, project AA2–17). M. K. is grateful for valuable discussions with Mark Friesen, Merritt P. Losert and Lars R. Schreiber.
@@ Appendix A Derivation of the Multi-Valley Coupled Envelope Equation Model
and the confinement potential $U\left(\mathbf{r}\right)$ described in Eq. (4). Here, $m_{0}$ is the vacuum electron mass and $V_{0}\left(\mathbf{r}\right)$ is the lattice-periodic potential of the crystal. The crystal potential is invariant under translations with lattice vectors $V_{0}\left(\mathbf{r}\right)\equiv V_{0}\left(\mathbf{r}+\mathbf{R}_{j}\right)$. We assume that $V_{0}$ describes solely the perfectly periodic Si crystal, whereas the effects of Ge in both the QW and the barrier are described by the heterostructure potential that is part of $U\left(\mathbf{r}\right)$, see Eq. (4). In this way, double-counting of Ge atoms is avoided and alloy disorder effects are separated from the idealized band structure computation.
@@ A.1 Lattice-Periodic Part
The lattice-periodic part of the coupled envelope equations is obtained from
@@ A.2 Heterostructure Part
In the following, the corresponding computation is carried out for the heterostructure part, which finally yields a coupling between the different valley-specific envelope wave functions. The ansatz (48) with (50) yields
@@
This expression is separated into the intravalley ($\nu=\nu^{\prime}$) and intervalley ($\nu\neq\nu^{\prime}$) contributions. Since the Bloch factors vary on an atomistic scale, i.e., much faster than the confinement potential and the envelope wave function, the intravalley part is obtained as [67]
@@ A.3 Coupled Envelopes
In the case of a single band (conduction band only, band indices $n$, $n^{\prime}$ are suppressed in the following), two valleys at $\mathbf{k}_{\nu}=\pm\mathbf{k}_{0}$, symmetry factors $\alpha_{\nu}=1/\sqrt{2}$ and identical effective masses $m_{n,\mathbf{k}_{\nu}}\equiv m$ and band edge energies $E_{n,\mathbf{k}_{\nu}}\equiv E_{c}$, we obtain
@@ Appendix B Evaluation of Valley Splitting Components and Band Structure Coefficients
In the small strain limit, the reciprocal lattice vectors are written as $\mathbf{G}_{n}=\sum_{j=1}^{3}n_{j}\left(I-\varepsilon\right)\mathbf{b}_{j}$, where the $\mathbf{b}_{j}$ are the primitive reciprocal lattice vectors of the relaxed fcc lattice [56]
@@
These selection rules greatly simplify the evaluation of the summation in $\Delta_{\mathrm{det}}^{l}$, which (after renaming the remaining summation index to $n$) yields the compact result
@@ C.1 Intervalley Coupling Parameter
where we have introduced the shorthand notation
@@
$=\mathrm{e}^{-\frac{1}{4}\left|s\right|^{2}\Gamma-\frac{1}{4}\mathrm{Re}\left(s^{*2}C\right)}.$
This is the characteristic function of a scalar complex-valued normal distribution with covariance $\Gamma$ and relation parameter (or pseudo-covariance) $C$. This is essentially the statement of the central limit theorem for complex-valued random variables. Hence